FIND THE EXACT VALUES WHEN ANGLE = 11pi/6 . DRAW ANGLE ON A UNIT CIRCLE?
Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(10682)


Hi
FIND THE EXACT VALUES WHEN ANGLE = 11pi/6 Angle = 330°

Answer by stanbon(57967)

FIND THE EXACT VALUES WHEN ANGLE = 11pi/6 . DRAW ANGLE ON A UNIT CIRCLE?
---
(11/6)pi is the same as -(1/6)pi or -30 degrees
---
If you draw it on the unit circle,
the radius is "1"
the y value is "-1/2"
the x value is "(1/2)sqrt(3)"
-------------------------------------
sin = -1/2
cos = sqrt(3)/2
tan = -sin/con = -1/sqrt(2)
-----
csc = -2
sec = 2/sqrt(2)
ctn = -sqrt(2)
========================
Cheers,
Stan H.